Write down as many correct, distinct, and relevant facts you can about the following prompt:

You wish to conduct a reaction in which sulfur dioxide (SO2) is produced. Your lab manual states that you need 0.05 cm3 of Sulfur-34. Due to sulfur being a solid at room temperature with a density of 2.07 g/cm3, you use a balance to weigh out your sample.

After the reaction is completed, you collect 2.2 x 10-6 grams of sulfur dioxide.

The statements must be relevant to the material that has been covered in this course. You can use outside information (such as reference material) but that will only count as one unique statement regardless of how much information you include that is from outside sources. For prompts that have you calculate a mass, solving for the molecular mass and the number of moles for a compound, count as only one distinct fact. Distinct facts must be independent and unique from any other statements you have made and this means for a unique calculation it must be different from any other calculations you have performed.


1
Expert's answer
2021-02-05T04:07:34-0500

Volume = 0.05 cm3 .

Density = mass / volume .

2.07 g / cm3 = mass / 0.05

Mass = ( 2.07 × 0.05 ) g = 0.1035 g .



But given mass ( experimental ) = 2.2 × 10-6 g.


% yield = 2.2 × 10-6 × 100 / 0.1035